# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ V-C3 Resource Allocation
Spectral Efficiency |
Energy Efficiency |
SL | UL | RL |
RL Model |
DL Model |
[114] |
AC |
[223] |
Q-learning |
[224] |
DQN |
FCNN |
[225] |
DQN |
CNN |
[226] |
DQN |
FCNN |
[227] |
MARL |
[228] |
DQN |
Ensembles of FCNN |
[229] |
FCNN |
[230] |
FCNN |
@@
Power and spectrum are the two fundamental resources for any type of wireless network, and NTN is also not an exception. The spectrum allocation is typically performed by the assignment of carriers with equal width from the allocated spectrum for that service. Hence, the number of assigned carriers and their positions are optimized to achieve good signal quality with the minimum resources. Often, the carrier assignment is achieved by the orthogonal splitting of the spectrum resources, which is also known as frequency reuse. However, the strict orthogonality of the frequency bands cannot be always achieved to achieve better spectral efficiency. In case of lack of orthogonality of spectrum resources used by different transceivers can also introduce CCI. The interfering signal can be effectively suppressed by increasing transmission power for the original signals. However, as power is also a scarce resource, we cannot increase the transmission power indefinitely and increasing transmission power will result in a decrease in energy efficiency. For better resource utilization, a more robust radio resource management needs to be designed by controlling both power and spectrum resources [231].
@@
Generally, an optimization framework can be considered to optimize the system performance with bandwidth and power constraints. In most cases, such optimization problems are non-linear and non-convex due to objective function nonlinearity and complex constraints involving Signal to Interference and Noise Ratio (SINR) [232]. Furthermore, the carrier assignment indicator variables result in a mixed-integer programming problem [232]. Hence, no optimal solution can be determined using the known methods of convex optimization with low computation complexity. Instead, suboptimal and metaheuristic approaches are proposed, which tackle parts of the problem separately and then iteratively tune the parameters [232]. Different suboptimal approaches are adopted to optimize resource allocations [233, 234, 235, 236, 237] for satellite systems. However to reduce computation complexity several heuristic [238] and metaheuristic approaches like GA [239], PSO [240] are explored to reach the suboptimal solutions within a shorter computation time.
@@
To tackle this resource allocation issue in real satellite networks in a practical manner, ML approaches are being started to be adopted by the research community. A DL framework is combined with conventional optimization algorithms to overcome the computation complexity issue of the conventional approach in [230, 229] by reducing the feature space. A model-free DRL framework is adopted for power allocation of high throughput satellites in [226]. A Q-learning-based long-term capacity allocation algorithm in an RL framework is introduced for a heterogeneous satellite network in [223]. In [114], an Actor-Critic and Critic Only based RL framework is considered for optimal resource allocation for LEO satellite networks. Different advanced RL frameworks like DRL [224, 241, 225], Multi-objective DRL [228] and MADRL [227] are also proposed to solve the resource allocation issue for satellites. In table VII, we summarize the AI approaches for resource allocation in TNTN. As both power and spectrum are equally important and scarce resources for NTNs, new DL architectures need to be explored to jointly allocate these resources in an efficient manner for NTNs.
@@ V-C4 Network Slicing
In a general network slicing framework, a composite utility function consisting of different network performance characteristics like average throughput and other costs like slice reconfiguration cost, resource reservation cost, etc. is formulated as an objective function that needs to be minimized. The constraints are generally the minimum service level To ensure real-time implementation, simple heuristic approaches are tested on real platforms [242, 243, 244]. requirements depending on the type of services for particular slices. In [242], an extensible 5G network slicing framework in conjunction with satellite networks is discussed to facilitate the integration of satellite services into 5G. In [245], a multi-objective optimization problem comprising latency, computational, and power requirements in an edge-computing scenario is formulated to find suitable slice scheduling strategies based on numerical methods. However, these approaches do not guarantee optimal performance guarantee. To tackle this issue, different AI-based approaches are explored as it is done in the case of traditional 5G terrestrial networks.
In [246], RL-based network slicing frameworks for satellite-integrated future 6G networks are discussed along with experimental results for simple networks. In [247], AI-based network slicing for space-air-ground integrated vehicular networks is discussed from the perspective of slice creation, user association, and resource scheduling. In [248, 249], satellite-terrestrial network slice resource allocation frameworks utilizing network function virtualization are presented which can be leveraged for applying advanced AI-based methods. In [250], FCNNs are used to train a suitable set of network parameters that can produce latency similar to a non-linear optimizer-based network slicer. In [251], a general Radio Access Network (RAN) slicing problem is considered where the objective function is a weighted function of bandwidth and spectrum consumption satisfying QoS and inter-slice isolation constraints. In a simple 2-slice satellite-terrestrial integrated network, different DL architectures are tested. In [252], an ML approach similar to the meta-heuristic ACO approach is considered to realize network slicing in a TNTN environment. An air-ground integrated network is considered in a DRL framework in [253], later solved by the DDPG algorithm. Here both the actor and critic networks are FCNNs consisting of four layers. Distributed learning architectures can be potentially explored in future works for real network implementations.
@@ V-C5 Handover Optimization
Signal Quality |
Network Load |
Service Time |
Tx Delay |
SL | UL | RL |
RL Model |
DL Model |
[254] |
Q-Learning |
[255] |
Online Q-Learning |
[256] |
Multi-Agent Q-Learning |
[257] |
DQN |
CNN |
[258] |
CNN |
[259] |
Auction based approach (Game theory) |
FCNN |
[260] |
DDQN |
FCNN |
[261] |
Successive DQN |
FCNN |
In order to maintain an orbital path around the Earth at a lower altitude, an LEO satellite needs to move at a much higher velocity (around km/s) compared to a GEO satellite. So these satellites orbit around the Earth typically within hours [20]. Due to the smaller orbital period, any LEO satellite remains visible to a ground UE for only several minutes which poses a great challenge for integrating these satellites into the traditional terrestrial networks. The UE needs to undergo multiple handovers within a short span of time interval regardless of its mobility status for seamless continuation of the data sessions [14]. This frequent handover phenomenon in LEO satellite networks creates a lot of overhead in communication channels and results in overall degradation in network performances. Moreover, due to lower altitudes, the coverage area of an LEO satellite is much smaller compared to a GEO satellite. Typically a large number of LEO satellites are needed to maintain global coverage across the Earth with complex constellations. In the case of an ultra-dense constellation of LEO satellites (like Starlink), each UE is generally covered by multiple satellites, so the UE can choose the best one from the list of suitable candidate LEO satellites. This problem can be potentially solved in an optimization framework jointly considering different handover decision criteria.
@@
Different simple greedy strategies, like Maximum Service Time (MST), Maximum Signal Quality (MSQ), or Minimum Network Load (MNL) [262] are adopted to solve the problem in a simple heuristic manner but none of these approaches provide the optimal solution. The satellite handover scenario can be also modeled as a directed graph between different satellites for a single user where the weights can be set by different handover criteria like Quality of Service (QoS), service time, etc. [263, 264]. A bipartite graph matching problem between the satellites and the users [265] is also considered in the literature to provide the optimal handover decision for satellites. In addition, a network flow-based cost minimization approach is considered in [266] by weighting each edge as the QoS perceived by the user. A handover strategy based on a potential game in a bipartite graph is considered in [267]. Different heuristic algorithms are also proposed to solve the problem [268, 269, 270]. A dynamic optimization problem is considered to be solved based on forecasting in [271]. Channel reservation is also associated to design an efficient handover algorithm while balancing the load for satellites in [272].
@@
An RL framework can be naturally adopted for solving this problem considering the handover criteria as states and UEs as agents who act by selecting a suitable LEO satellite and collecting a reward based on the network performances. In [254], only the overall signal quality of the network is maximized using the RL approach without considering any other criteria. In [257, 255], a multi-objective optimization problem considering satellite load and signal quality constraints is solved using the DRL approach. In real networks, we have a large number of UEs; the handover decision for one UE can affect another UE, so the handover problem needs to be solved in a cooperative manner. In [256], a MARL framework is considered where multiple UEs cooperatively optimize the number of handovers in the whole network considering different handover criteria. In [258], using graph matching, a database of optimum handover decisions in satellite networks is produced and later it is used to predict handover decisions using a CNN model. Advanced DL architectures like Auction based DL [259], DDQN [260], Successive DQN [261], etc. are also considered to provide optimal handover decisions. In table VIII, we summarize the AI approaches for handover optimization in NTN involving LEO satellites. However, as all the system models consider the agents located at the UE side, it does not comply with the current standards where the handover decision is generally controlled by the BSs (satellites in this case). Furthermore, the distributed multi-agent learning architectures give rise to stability issues in real implementations. The handover criteria also need to be carefully investigated to provide the agents with the necessary information to learn the mobility behavior of the environment. These issues need to be resolved in an efficient manner for future research work in this domain.
@@ V-C6 Multiple Access
The optimum power allocation problem in NOMA for NTNs can be formulated as a non-convex problem which is often difficult to solve using conventional approaches. In[273], a long-term power allocation scheme for NOMA in satellite-IoT networks is solved by deriving the optimal decoding order leveraging DL techniques. A DQN-based DRL approach is investigated in [274] for optimum power allocation in satellite-IoT networks under different channel conditions and delay-QoS requirements of NOMA users. In another study as discussed in [275], the non-convex problem involving integer variables is later reformulated as a mixed-integer convex problem which was later solved by two DL techniques instead of conventional iterative solutions. Some studies also focus on the non-convex user selection problem for given power allocations based on the CSI feedback. Such a study, [276] used DQN to find out the suitable user pairing for delay-limited NOMA-based satellite networks considering the channel conditions and delay constraints as states. k-means UL approach is also considered to find out the pair of terrestrial users to be simultaneously served by space and aerial BSs adopting NOMA in [277]. Q-learning is adopted in [278] to allocate the time slots and communications channels for IoT-satellite terrestrial relay networks.
@@
Another significant multiple access method, RSMA, is also explored in the context of satellite networks to enhance spectral efficiency. In RSMA, user messages are partitioned into two segments: common and private. The common signals are collectively encoded and merged into a unified data stream intended for all users, while each user linearly precodes their private messages. On the receiver side, the common component is extracted while treating the private signals as noise, employing the SIC technique. Subsequently, each user extracts their respective private signals. This provides the users with another way to share the same resource blocks with an increase in spectral efficiency. Generally maximizing the sum rate for both parts is a complicated non-convex problem and can be solved by Weighted MMSE (WMMSE) problem which is difficult to implement in practical hardware. A successive convex approximation as well as KarushKuhnTucker (KKT) conditions are used to calculate the transmit power in RSMA power for different beams in satellite networks in [279]. However, DL techniques can be extremely useful in modeling the solution framework with low complexity as shown in [280, 281]. Here a deep unfolding technique is used to implement the WMMSE algorithm using a deep NN and momentum-accelerated Projection Gradient Descent algorithm. A DRL framework using Proximal Policy Optimization is used to maximize the sum rate in [282]. Here each BS works as an agent, the channel state information i.e. SINR of the private and common messages is used as the states, the action is to find the suitable power allocation whereas the reward is the achieved sum rate.
@@ V-D1 Computation Offloading
Energy Consumption |
Tx Delay |
Computational Resources |
SL | UL | RL |
RL Model |
DL Model |
[283] |
DDPG |
FCNN |
[284] |
Model-Free |
FCNN |
[285] |
Value Iteration, DQN, DDQN |
FCNN |
[285] |
Dueling DDQN |
FCNN |
[286] |
MARL |
[287] |
DQN based MARL |
FCNN |
[288] |
LSTM |
[289] |
DQN |
FCNN |
[290] |
Distributed FCNN (For solving Optimization) |
[291] |
Distributed FCNN (For solving Optimization) |
@@
One of the most important applications of satellite-terrestrial integrated networks is enhancing the computation capabilities of existing terrestrial network architectures leveraging satellites. With traditional terrestrial networks, supporting a diverse set of new applications like AR, VR, etc. with high data processing and extremely low latency requirement can get very challenging. Generally, terrestrial BSs are deployed sparsely due to high infrastructure and maintenance costs. Due to resource constraints, in case of the high demand for data processing for these types of applications, the BSs need to offload the computation tasks to the terrestrial cloud via the satellites [292]. However, due to longer propagation delay, the latency requirements set by the applications are difficult to be met [293]. Nevertheless, due to the emergence of LEO satellites with comparatively low propagation delay, the overall delay is considerably reduced. Also instead of acting as relays, the satellite can now also do the processing works acting as edge-servers. So we can consider a three-level hierarchical architecture comprising of ground UEs connected to terrestrial BSs, LEO satellites, and terrestrial cloud as shown in Figure 17 where the terrestrial BSs can offload the computational tasks to LEO satellites and to terrestrial clouds via the LEO satellites.
@@
The main challenges in task offloading problems lie in meeting the delay constraints for low-latency applications while minimizing the energy consumption for the satellites. So this can be formulated as an optimization problem to come up with an efficient offloading approach for integrated TNTN architecture. Such an optimization problem is solved using different conventional approaches like 3D hypergraph matching [294], game theory [295], stochastic approach [296], efficient algorithms [297] in the existing literature. In [298], a joint optimization framework comprising task offloading and resource allocation is also considered in an integrated satellite-terrestrial environment. Although these algorithms work well in theory for particular scenarios, in real networks, the feasibility of these algorithms is compromised due to different issues. Some of these works do not consider the cooperation among terrestrial cloud and LEO satellite servers which result in sub-optimal approaches [294, 295]. Also, these approaches are some predefined models highly dependent on different network states causing a large overhead in networks. Moreover, they usually converge to the solutions after a large number of iterations causing high computational complexity.
@@
To tackle these issues, different ML approaches are proposed in the literature to solve task offloading problems. In [284], a DRL-based task offloading framework dependent on channel state information is proposed. A similar DRL-based framework is also considered in [299] with additional consideration of the dynamic queue condition in satellites. In [285], both DQN and DDQN are explored to solve the task offloading problem in a decentralized manner. DDPG algorithm is considered to solve the optimization problem in a DQN framework in [283] while taking the potential security issues into account. An LSTM model is used to solve the task offloading problem while considering channel conditions and energy dynamics in [288]. A DL-based caching strategy is considered in satellite edge networks in [289]. As we have multiple satellites in the real networks, to improve the overall system performances, different multi-agent architectures are considered both in a distributed [290] and cooperative environment [286, 287]. Distributed architectures for generating discrete offloading decisions in a supervised manner are also considered in [291]. In table IX, we summarize the AI approaches for task offloading in TNTN. As the delay constraints vary with network traffic types, the offloading decisions need to be derived taking network traffic types into account. Potential research works can show how computational offloading can be done for various network traffics and show superior network performances.

### v4

@@ V-C3 Resource Allocation
| Spectral Efficiency | Energy Efficiency | SL | UL | RL | RL Model | DL Model | ||
| [ntn-ra-leo-rl1] | AC | |||||||
| [ntn-ra-mul-rl2] | Q-learning | |||||||
| [ntn-ra-drl1] | DQN | FCNN | ||||||
| [ntn-ra-drl3] | DQN | CNN | ||||||
| [ntn-ra-drl4] | DQN | FCNN | ||||||
| [ntn-ra-madrl1] | MARL | |||||||
| [ntn-ra-modrl-ens] | DQN | Ensembles of FCNN | ||||||
| [ntn-ra-dl1] | FCNN | |||||||
| [ntn-ra-dl2] | FCNN | |||||||
@@
Power and spectrum are the two fundamental resources for any type of wireless network, and NTN is also not an exception. The spectrum allocation is typically performed by the assignment of carriers with equal width from the allocated spectrum for that service. Hence, the number of assigned carriers and their positions are optimized to achieve good signal quality with the minimum resources. Often, the carrier assignment is achieved by the orthogonal splitting of the spectrum resources, which is also known as frequency reuse. However, the strict orthogonality of the frequency bands cannot be always achieved to achieve better spectral efficiency. In case of lack of orthogonality of spectrum resources used by different transceivers can also introduce CCI. The interfering signal can be effectively suppressed by increasing transmission power for the original signals. However, as power is also a scarce resource, we cannot increase the transmission power indefinitely and increasing transmission power will result in a decrease in energy efficiency. For better resource utilization, a more robust radio resource management needs to be designed by controlling both power and spectrum resources [ntn-ra-gen2].
@@
Generally, an optimization framework can be considered to optimize the system performance with bandwidth and power constraints. In most cases, such optimization problems are non-linear and non-convex due to objective function nonlinearity and complex constraints involving Signal to Interference and Noise Ratio (SINR) [ntn-ra-gen1]. Furthermore, the carrier assignment indicator variables result in a mixed-integer programming problem [ntn-ra-gen1]. Hence, no optimal solution can be determined using the known methods of convex optimization with low computation complexity. Instead, suboptimal and metaheuristic approaches are proposed, which tackle parts of the problem separately and then iteratively tune the parameters [ntn-ra-gen1]. Different suboptimal approaches are adopted to optimize resource allocations [ntn-ra-opt1, ntn-ra-opt2, ntn-ra-opt3, ntn-ra-opt4, ntn-ra-opt5] for satellite systems. However to reduce computation complexity several heuristic [ntn-ra-heu1] and metaheuristic approaches like GA [ntn-ra-ga], PSO [ntn-ra-pso] are explored to reach the suboptimal solutions within a shorter computation time.
@@
To tackle this resource allocation issue in real satellite networks in a practical manner, ML approaches are being started to be adopted by the research community. A DL framework is combined with conventional optimization algorithms to overcome the computation complexity issue of the conventional approach in [ntn-ra-dl2, ntn-ra-dl1] by reducing the feature space. A model-free DRL framework is adopted for power allocation of high throughput satellites in [ntn-ra-drl4]. A Q-learning-based long-term capacity allocation algorithm in an RL framework is introduced for a heterogeneous satellite network in [ntn-ra-mul-rl2]. In [ntn-ra-leo-rl1], an Actor-Critic and Critic Only based RL framework is considered for optimal resource allocation for LEO satellite networks. Different advanced RL frameworks like DRL [ntn-ra-drl1, ntn-ra-drl2, ntn-ra-drl3], Multi-objective DRL [ntn-ra-modrl-ens] and MADRL [ntn-ra-madrl1] are also proposed to solve the resource allocation issue for satellites. In table VII, we summarize the AI approaches for resource allocation in TNTN. As both power and spectrum are equally important and scarce resources for NTNs, new DL architectures need to be explored to jointly allocate these resources in an efficient manner for NTNs.
@@ V-C4 Network Slicing
In a general network slicing framework, a composite utility function consisting of different network performance characteristics like average throughput and other costs like slice reconfiguration cost, resource reservation cost, etc. is formulated as an objective function that needs to be minimized. The constraints are generally the minimum service level To ensure real-time implementation, simple heuristic approaches are tested on real platforms [ns-gen1, ns-gen2, ns-gen3]. requirements depending on the type of services for particular slices. In [ns-gen1], an extensible 5G network slicing framework in conjunction with satellite networks is discussed to facilitate the integration of satellite services into 5G. In [ns-gen4], a multi-objective optimization problem comprising latency, computational, and power requirements in an edge-computing scenario is formulated to find suitable slice scheduling strategies based on numerical methods. However, these approaches do not guarantee optimal performance guarantee. To tackle this issue, different AI-based approaches are explored as it is done in the case of traditional 5G terrestrial networks.
In [ns-ai-tn], RL-based network slicing frameworks for satellite-integrated future 6G networks are discussed along with experimental results for simple networks. In [ns-ai-vntn], AI-based network slicing for space-air-ground integrated vehicular networks is discussed from the perspective of slice creation, user association, and resource scheduling. In [ns-vnf, ns-vnf2], satellite-terrestrial network slice resource allocation frameworks utilizing network function virtualization are presented which can be leveraged for applying advanced AI-based methods. In [ns-dl], FCNNs are used to train a suitable set of network parameters that can produce latency similar to a non-linear optimizer-based network slicer. In [ns-ai], a general Radio Access Network (RAN) slicing problem is considered where the objective function is a weighted function of bandwidth and spectrum consumption satisfying QoS and inter-slice isolation constraints. In a simple 2-slice satellite-terrestrial integrated network, different DL architectures are tested. In [ns-ai2], an ML approach similar to the meta-heuristic ACO approach is considered to realize network slicing in a TNTN environment. An air-ground integrated network is considered in a DRL framework in [ns-ddrl], later solved by the DDPG algorithm. Here both the actor and critic networks are FCNNs consisting of four layers. Distributed learning architectures can be potentially explored in future works for real network implementations.
@@ V-C5 Handover Optimization
| Signal Quality | Network Load | Service Time | Tx Delay | SL | UL | RL | RL Model | DL Model | ||
| [ho-sq] | Q-Learning | |||||||||
| [ho-qoe] | Online Q-Learning | |||||||||
| [ho-marl] | Multi-Agent Q-Learning | |||||||||
| [ho-drl] | DQN | CNN | ||||||||
| [ho-ai] | CNN | |||||||||
| [ho-auc-dl] | Auction based approach (Game theory) | FCNN | ||||||||
| [ho-ddqn] | DDQN | FCNN | ||||||||
| [ho-sdqn] | Successive DQN | FCNN | ||||||||
In order to maintain an orbital path around the Earth at a lower altitude, an LEO satellite needs to move at a much higher velocity (around km/s) compared to a GEO satellite. So these satellites orbit around the Earth typically within hours [3gpp1]. Due to the smaller orbital period, any LEO satellite remains visible to a ground UE for only several minutes which poses a great challenge for integrating these satellites into the traditional terrestrial networks. The UE needs to undergo multiple handovers within a short span of time interval regardless of its mobility status for seamless continuation of the data sessions [NTN_2]. This frequent handover phenomenon in LEO satellite networks creates a lot of overhead in communication channels and results in overall degradation in network performances. Moreover, due to lower altitudes, the coverage area of an LEO satellite is much smaller compared to a GEO satellite. Typically a large number of LEO satellites are needed to maintain global coverage across the Earth with complex constellations. In the case of an ultra-dense constellation of LEO satellites (like Starlink), each UE is generally covered by multiple satellites, so the UE can choose the best one from the list of suitable candidate LEO satellites. This problem can be potentially solved in an optimization framework jointly considering different handover decision criteria.
@@
Different simple greedy strategies, like Maximum Service Time (MST), Maximum Signal Quality (MSQ), or Minimum Network Load (MNL) [ho-gen] are adopted to solve the problem in a simple heuristic manner but none of these approaches provide the optimal solution. The satellite handover scenario can be also modeled as a directed graph between different satellites for a single user where the weights can be set by different handover criteria like Quality of Service (QoS), service time, etc. [ho-graph, ho-graph3]. A bipartite graph matching problem between the satellites and the users [ho-graph2] is also considered in the literature to provide the optimal handover decision for satellites. In addition, a network flow-based cost minimization approach is considered in [ho-nf] by weighting each edge as the QoS perceived by the user. A handover strategy based on a potential game in a bipartite graph is considered in [ho-game]. Different heuristic algorithms are also proposed to solve the problem [ho-t1, ho-t2, ho-t3]. A dynamic optimization problem is considered to be solved based on forecasting in [ho-for]. Channel reservation is also associated to design an efficient handover algorithm while balancing the load for satellites in [ho-chlo].
@@
An RL framework can be naturally adopted for solving this problem considering the handover criteria as states and UEs as agents who act by selecting a suitable LEO satellite and collecting a reward based on the network performances. In [ho-sq], only the overall signal quality of the network is maximized using the RL approach without considering any other criteria. In [ho-drl, ho-qoe], a multi-objective optimization problem considering satellite load and signal quality constraints is solved using the DRL approach. In real networks, we have a large number of UEs; the handover decision for one UE can affect another UE, so the handover problem needs to be solved in a cooperative manner. In [ho-marl], a MARL framework is considered where multiple UEs cooperatively optimize the number of handovers in the whole network considering different handover criteria. In [ho-ai], using graph matching, a database of optimum handover decisions in satellite networks is produced and later it is used to predict handover decisions using a CNN model. Advanced DL architectures like Auction based DL [ho-auc-dl], DDQN [ho-ddqn], Successive DQN [ho-sdqn], etc. are also considered to provide optimal handover decisions. In table VIII, we summarize the AI approaches for handover optimization in NTN involving LEO satellites. However, as all the system models consider the agents located at the UE side, it does not comply with the current standards where the handover decision is generally controlled by the BSs (satellites in this case). Furthermore, the distributed multi-agent learning architectures give rise to stability issues in real implementations. The handover criteria also need to be carefully investigated to provide the agents with the necessary information to learn the mobility behavior of the environment. These issues need to be resolved in an efficient manner for future research work in this domain.
@@ V-C6 Multiple Access
The optimum power allocation problem in NOMA for NTNs can be formulated as a non-convex problem which is often difficult to solve using conventional approaches. In[ntn-noma1], a long-term power allocation scheme for NOMA in satellite-IoT networks is solved by deriving the optimal decoding order leveraging DL techniques. A DQN-based DRL approach is investigated in [ntn-noma2] for optimum power allocation in satellite-IoT networks under different channel conditions and delay-QoS requirements of NOMA users. In another study as discussed in [ntn-noma3], the non-convex problem involving integer variables is later reformulated as a mixed-integer convex problem which was later solved by two DL techniques instead of conventional iterative solutions. Some studies also focus on the non-convex user selection problem for given power allocations based on the CSI feedback. Such a study, [ntn-noma4] used DQN to find out the suitable user pairing for delay-limited NOMA-based satellite networks considering the channel conditions and delay constraints as states. k-means UL approach is also considered to find out the pair of terrestrial users to be simultaneously served by space and aerial BSs adopting NOMA in [ntn-noma5]. Q-learning is adopted in [ntn-noma6] to allocate the time slots and communications channels for IoT-satellite terrestrial relay networks.
@@
Another significant multiple access method, RSMA, is also explored in the context of satellite networks to enhance spectral efficiency. In RSMA, user messages are partitioned into two segments: common and private. The common signals are collectively encoded and merged into a unified data stream intended for all users, while each user linearly precodes their private messages. On the receiver side, the common component is extracted while treating the private signals as noise, employing the SIC technique. Subsequently, each user extracts their respective private signals. This provides the users with another way to share the same resource blocks with an increase in spectral efficiency. Generally maximizing the sum rate for both parts is a complicated non-convex problem and can be solved by Weighted MMSE (WMMSE) problem which is difficult to implement in practical hardware. A successive convex approximation as well as KarushKuhnTucker (KKT) conditions are used to calculate the transmit power in RSMA power for different beams in satellite networks in [ntn-rsma4]. However, DL techniques can be extremely useful in modeling the solution framework with low complexity as shown in [ntn-rsma3, ntn-rsma5]. Here a deep unfolding technique is used to implement the WMMSE algorithm using a deep NN and momentum-accelerated Projection Gradient Descent algorithm. A DRL framework using Proximal Policy Optimization is used to maximize the sum rate in [ntn-rsma2]. Here each BS works as an agent, the channel state information i.e. SINR of the private and common messages is used as the states, the action is to find the suitable power allocation whereas the reward is the achieved sum rate.
@@ V-D1 Computation Offloading
| Energy Consumption | Tx Delay | Computational Resources | SL | UL | RL | RL Model | DL Model | ||
| [to-rl] | DDPG | FCNN | |||||||
| [to-drl1] | Model-Free | FCNN | |||||||
| [to-drl2] | Value Iteration, DQN, DDQN | FCNN | |||||||
| [to-drl2] | Dueling DDQN | FCNN | |||||||
| [to-marl] | MARL | ||||||||
| [to-madrl] | DQN based MARL | FCNN | |||||||
| [to-dl1] | LSTM | ||||||||
| [to-dl2] | DQN | FCNN | |||||||
| [to-ddl1] | Distributed FCNN (For solving Optimization) | ||||||||
| [to-ddl2] | Distributed FCNN (For solving Optimization) | ||||||||
@@
One of the most important applications of satellite-terrestrial integrated networks is enhancing the computation capabilities of existing terrestrial network architectures leveraging satellites. With traditional terrestrial networks, supporting a diverse set of new applications like AR, VR, etc. with high data processing and extremely low latency requirement can get very challenging. Generally, terrestrial BSs are deployed sparsely due to high infrastructure and maintenance costs. Due to resource constraints, in case of the high demand for data processing for these types of applications, the BSs need to offload the computation tasks to the terrestrial cloud via the satellites [to-ov1]. However, due to longer propagation delay, the latency requirements set by the applications are difficult to be met [to-ov2]. Nevertheless, due to the emergence of LEO satellites with comparatively low propagation delay, the overall delay is considerably reduced. Also instead of acting as relays, the satellite can now also do the processing works acting as edge-servers. So we can consider a three-level hierarchical architecture comprising of ground UEs connected to terrestrial BSs, LEO satellites, and terrestrial cloud as shown in Figure 17 where the terrestrial BSs can offload the computational tasks to LEO satellites and to terrestrial clouds via the LEO satellites.
@@
The main challenges in task offloading problems lie in meeting the delay constraints for low-latency applications while minimizing the energy consumption for the satellites. So this can be formulated as an optimization problem to come up with an efficient offloading approach for integrated TNTN architecture. Such an optimization problem is solved using different conventional approaches like 3D hypergraph matching [to-gen1], game theory [to-game], stochastic approach [to-stoch], efficient algorithms [to-gen2] in the existing literature. In [to-joint], a joint optimization framework comprising task offloading and resource allocation is also considered in an integrated satellite-terrestrial environment. Although these algorithms work well in theory for particular scenarios, in real networks, the feasibility of these algorithms is compromised due to different issues. Some of these works do not consider the cooperation among terrestrial cloud and LEO satellite servers which result in sub-optimal approaches [to-gen1, to-game]. Also, these approaches are some predefined models highly dependent on different network states causing a large overhead in networks. Moreover, they usually converge to the solutions after a large number of iterations causing high computational complexity.
@@
To tackle these issues, different ML approaches are proposed in the literature to solve task offloading problems. In [to-drl1], a DRL-based task offloading framework dependent on channel state information is proposed. A similar DRL-based framework is also considered in [to-drl3] with additional consideration of the dynamic queue condition in satellites. In [to-drl2], both DQN and DDQN are explored to solve the task offloading problem in a decentralized manner. DDPG algorithm is considered to solve the optimization problem in a DQN framework in [to-rl] while taking the potential security issues into account. An LSTM model is used to solve the task offloading problem while considering channel conditions and energy dynamics in [to-dl1]. A DL-based caching strategy is considered in satellite edge networks in [to-dl2]. As we have multiple satellites in the real networks, to improve the overall system performances, different multi-agent architectures are considered both in a distributed [to-ddl1] and cooperative environment [to-marl, to-madrl]. Distributed architectures for generating discrete offloading decisions in a supervised manner are also considered in [to-ddl2]. In table IX, we summarize the AI approaches for task offloading in TNTN. As the delay constraints vary with network traffic types, the offloading decisions need to be derived taking network traffic types into account. Potential research works can show how computational offloading can be done for various network traffics and show superior network performances.